OpenStudy (anonymous):

Solve the system of equations: 3x + 8y = 72 -5x - 2y = -52

OpenStudy (souvik):

multiply 2nd equation with 4 and add them

OpenStudy (anonymous):

4? How do you even get 4?

OpenStudy (souvik):

if you do you it...y would be eliminated

OpenStudy (anonymous):

That doesn't answer my question.. Where do you even get a 4 out of this equation?

OpenStudy (anonymous):

Ho did you figure t that you need to multiply it by 4?

OpenStudy (texaschic101):

The elimination method works by eliminating the y's or the x's. To do this, sometimes you have to multiply one or sometimes both of the equations by a number for this to occur. souvik is right in telling you to multiply the 2nd equation by 4 because if you do that, you will find that the y's will cancel each other out. Try it. What do you get when you multiply the 2nd equation by 4 ?

OpenStudy (anonymous):

-20x-8y=-208

OpenStudy (texaschic101):

Let me explain more....if you multiply the 2nd equation by 4 you will get -20x - 8y = - 208. Now, when you add the equations, you will see that the 1st equation has a 8y and the 2nd equation has a -8y, and they cancel out. Do you understand ?

OpenStudy (anonymous):

Yes

OpenStudy (texaschic101):

3x + 8y = 72 -5x - 2y = -52 -->(4)-5x - 2y = -52 -------------- 3x + 8y = 72 -20x -8y = - 208 ---------------add -17x = - 136 x = 8 now sub 8 in for x in either of the original equations to get y. Does this make sense to you ?

OpenStudy (texaschic101):

This is the easiest way, however, you could have multiplied the 1st equation by 5 and the 2nd equation by 3, and that would make the x's cancel out. Either way, you should come up with the same answer.

OpenStudy (anonymous):

Yes, I get it now..

OpenStudy (texaschic101):

good to hear :)

OpenStudy (anonymous):

(8, 6) ?

OpenStudy (texaschic101):

Let me check...hold on

OpenStudy (marissalovescats):

x=8 and y=6 using the elimination method as they are describing.

OpenStudy (souvik):

right...!

OpenStudy (texaschic101):

x = 8 now sub 8 in for x 3x + 8y = 72 3(8) + 8y = 72 24 + 8y = 72 8y = 72 - 24 8y = 48 y = 6 you can check your answers by subbing in your info in either of the original equations... -5x - 2y = -52 -5(8) -2(6) = -52 -40 - 12 = - 52 -52 = -52 (correct) If they come out even, then it is correct, if not, then a mistake was made or their is no solution. Yes....answer is (8,6)

OpenStudy (anonymous):

Yay, thank you.! c:

OpenStudy (texaschic101):

anytime :)

OpenStudy (marissalovescats):

Yay math!